OpenStudy (toxicsugar22):

sketch the triangle with vertics A(-5,12), B(4,-2) and C(1,-6) Find the slope of the side AC. Find the slope of the altitude form point B to side AC fIND an equation for the line that includes the altitude from point B to side AC

OpenStudy (anonymous):

Do you recall the formula for the slope of a line given 2 points on the line?

OpenStudy (toxicsugar22):

yes

OpenStudy (anonymous):

So use it to find the slope of the side AC.

OpenStudy (anonymous):

Once you have that you can find the perpendicular slope by taking the negative reciprocal

OpenStudy (anonymous):

which is the slope of the altitude from B to the side AC. Then with that slope you can write the equation of the line with that slope that intersects B.

OpenStudy (toxicsugar22):

im guessing -3

OpenStudy (toxicsugar22):

is that right

OpenStudy (anonymous):

-3 is the slope of AC

OpenStudy (toxicsugar22):

ok

OpenStudy (anonymous):

So what is the slope perpendicular to that?

OpenStudy (toxicsugar22):

so what would be the equatin

OpenStudy (toxicsugar22):

1/-3

OpenStudy (toxicsugar22):

right

OpenStudy (anonymous):

No, negative reciprocal is perpendicular.

OpenStudy (anonymous):

so 1/3

OpenStudy (anonymous):

So now, you have a point B and the slope. Do you know the formula for the equation of the line given a point and a slope?

OpenStudy (toxicsugar22):

y=1/3x+2

OpenStudy (toxicsugar22):

is that right

OpenStudy (toxicsugar22):

is thais right

OpenStudy (anonymous):

I don't think so. It should be \(y+2 = \frac{1}{3}(x-4)\)

OpenStudy (toxicsugar22):

i think we put it in slope intecept form

OpenStudy (anonymous):

sure, but I don't think if you do you'll get the answer you had before

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

vut can u help me get me put that in slope inetcept form

OpenStudy (anonymous):

Well start by distributing the 1/3, then subtract 2 from both sides

OpenStudy (toxicsugar22):

y+2=1/3(x-4) can u help me put tis in slpo intecept orm

OpenStudy (toxicsugar22):

can u help me

OpenStudy (anonymous):

did you distribute the 1/3?

OpenStudy (toxicsugar22):

how do u do -4/3-2

OpenStudy (anonymous):

convert -2 into an improper fraction (with a 3 in the denominator)

OpenStudy (toxicsugar22):

i got y+2=1/3x-4/3

OpenStudy (toxicsugar22):

is that right then i subract 2 from both sides

OpenStudy (anonymous):

yes

OpenStudy (toxicsugar22):

then i get y=1/3x-3.3

OpenStudy (toxicsugar22):

is this write

OpenStudy (anonymous):

-3.33333 = -10/3

OpenStudy (toxicsugar22):

thanks you so much

OpenStudy (toxicsugar22):

The center of a circle is the point C(2,4), and P(-1,6) is a point on the circle. Find the equation of the line tangent to the circle at the point P.

OpenStudy (toxicsugar22):

can u help with this

OpenStudy (anonymous):

umm.. it's pretty nearly the same problem, just with different numbers

OpenStudy (toxicsugar22):

ok

OpenStudy (toxicsugar22):

but can u help me

OpenStudy (anonymous):

draw the picture of what they're describing, and recall what we did in this last problem and see if you can figure out how the problems are similar.

OpenStudy (toxicsugar22):

i tried please can u help me

OpenStudy (anonymous):

what did you try?

OpenStudy (toxicsugar22):

i tried doing it but idid not get it

OpenStudy (toxicsugar22):

doing the pomble

OpenStudy (toxicsugar22):

can u help me

OpenStudy (anonymous):

what did you _do_ when you tried?

OpenStudy (toxicsugar22):

idid (2,4) and (- 1and 6) and added the x and y

OpenStudy (toxicsugar22):

please can u help, I will be greatful to you

OpenStudy (anonymous):

what were you trying to find when you did that?

OpenStudy (toxicsugar22):

idk i thought that was how u do that

OpenStudy (anonymous):

do what?

OpenStudy (toxicsugar22):

the question

OpenStudy (anonymous):

Are you trying to find something?

OpenStudy (toxicsugar22):

yes the equation

OpenStudy (anonymous):

the equation of what?

OpenStudy (toxicsugar22):

The center of a circle is the point C(2,4), and P(-1,6) is a point on the circle. Find the equation of the line tangent to the circle at the point P

OpenStudy (anonymous):

I know what the problem says, I'm asking you what you think. You do think right? ;)

OpenStudy (toxicsugar22):

i am finding it difficult please can u help me

OpenStudy (anonymous):

What are you trying to find here?

OpenStudy (toxicsugar22):

the equation i guess

OpenStudy (anonymous):

The equation of what?

OpenStudy (toxicsugar22):

the slope

OpenStudy (toxicsugar22):

the slope

OpenStudy (anonymous):

"Find the equation of the line tangent to the circle at the point P"

OpenStudy (anonymous):

What does that sentence mean?

OpenStudy (toxicsugar22):

i think it means that the line is touching line p

OpenStudy (toxicsugar22):

is that right

OpenStudy (toxicsugar22):

is this righ t

OpenStudy (anonymous):

The line is touching the point P that's on the circle, but what else do you know about it?

OpenStudy (toxicsugar22):

idk

OpenStudy (anonymous):

well, specifically how does it relate to the radius of the circle through the point P?

OpenStudy (toxicsugar22):

is the slope -3/2

OpenStudy (toxicsugar22):

is it right

OpenStudy (anonymous):

that is the slope of the radius (line between the center and P). The trick is to realize that the slope of the tangent will be perpendicular to the radius.

OpenStudy (toxicsugar22):

so it will be 2/3

OpenStudy (toxicsugar22):

is that right

OpenStudy (anonymous):

that is the slope. now use the point p to find the equation like we did before

OpenStudy (toxicsugar22):

so we do (-1,-6)

OpenStudy (toxicsugar22):

so it would be y+6=2/3(x+1)

OpenStudy (toxicsugar22):

is that right

OpenStudy (anonymous):

I thought P was (-1,6), so it would be y-6 = 2/3(x+1)

OpenStudy (toxicsugar22):

oh yeah sorry your right

OpenStudy (toxicsugar22):

if i put it in slope intecept orm it would be y=2/3x+6.67

OpenStudy (toxicsugar22):

is this right

OpenStudy (anonymous):

yeah

OpenStudy (toxicsugar22):

but is the 6.67 right or can i put it in a fraction

OpenStudy (anonymous):

you can make it a fraction

OpenStudy (toxicsugar22):

how can u show me

OpenStudy (anonymous):

How many 3rds are there in 6 wholes?

OpenStudy (toxicsugar22):

so id it 20/3

OpenStudy (toxicsugar22):

wait i think we messed up it would be y+6=2/3(X-1)

OpenStudy (toxicsugar22):

i think that is right

OpenStudy (toxicsugar22):

i think that is right

OpenStudy (anonymous):

If the point P is (-1,6) then the equation would be y-6=2/3(x+1) The formula for point-slope is: \[y-y_1 = m(x-x_1)\] Where \(x_1,y_1\) are the x,y values of your point

OpenStudy (toxicsugar22):

so would it be y=2/3x+3/20